OpenStudy (meganamerald):

How do I find the absolute value of 6-16i ? Imaginary numbers are awful

OpenStudy (sidsiddhartha):

HINT: absolute value of a complex quantity (A+iB) is \[A=\sqrt{A^2+B^2}\]

OpenStudy (sidsiddhartha):

\[Z=\sqrt{A^2+B^2}\]

OpenStudy (anonymous):

multiply by the congugate then take the square root to find the modulus... |6-16i| = sqrt{(6-16i)(6+16i)}

OpenStudy (anonymous):

if \(z=a+bi\) then \(|z|=\sqrt{zz^*}=\sqrt{(a+bi)(a-bi)}=\sqrt{a^2+b^2}\)

OpenStudy (meganamerald):

\[\sqrt{36+256}= \sqrt{292}\]

OpenStudy (meganamerald):

I dont understand what to do after that

OpenStudy (sidsiddhartha):

yes thats correct thats the absolute value of the complex quantity

OpenStudy (meganamerald):

Thats it?

OpenStudy (sidsiddhartha):

yup ,u could simplify it a little bit \[\sqrt{292}=17.08\]

OpenStudy (meganamerald):

Awesome thanks :)

OpenStudy (sidsiddhartha):

yw!!!